# CBQ: Cross-Block Quantization for Large Language Models

### v4

@@ Abstract
Post-training quantization (PTQ) has played a key role in compressing large language models (LLMs) with ultra-low costs. However, existing PTQ methods only focus on handling the outliers within one layer or one block, which ignores the dependency of blocks and leads to severe performance degradation in low-bit settings. In this paper, we propose CBQ, a cross-block reconstruction-based PTQ method for LLMs. CBQ employs a cross-block dependency using a homologous reconstruction scheme, establishing long-range dependencies across multiple blocks to minimize error accumulation. Furthermore, CBQ incorporates a coarse-to-fine preprocessing (CFP) strategy for suppressing weight and activation outliers, coupled with an adaptive LoRA-Rounding technique for precise weight quantization. These innovations enable CBQ to not only handle extreme outliers effectively but also improve overall quantization accuracy. Extensive experiments show that CBQ achieves superior low-bit quantization (W4A4, W4A8, W2A16) and outperforms existing state-of-the-art methods across various LLMs and datasets. Notably, CBQ quantizes the 4-bit LLAMA1-65B model within only 4.3 hours on a single GPU, achieving a commendable tradeoff between performance and quantization efficiency.
Machine Learning, ICML
(a) Layer-by-layer quantization (Xiao et al., 2022)
(b) Block-by-block quantization (Shao et al., 2023)
(c) Cross-block quantization (Ours)
Figure 1: Comparison of different optimization manners in post-training quantization for large language models. Orange and blue dashed boxes indicate the order of quantization optimization.
@@
1 Introduction
Large language models (LLMs) (Wei et al., 2022a; Radford et al., ; Zhang et al., ; Brown et al., 2020), have sparked immense academic and industrial interest owing to their remarkable performance in handling complex natural languages tasks (Hendrycks et al., 2020b; Bisk et al., 2020b), like language generation, translation, question answering, and text summarization etc. However, the sheer size of these models, often comprising billion-level parameters, demands significant computational resources for inference and deployment. This problem necessitates the employment of model compression techniques such as network pruning (Han et al., 2015), knowledge distillation (Hinton et al., 2015), network quantization (Choi et al., 2018; Frantar et al., 2022b), and neural architecture search (Zoph & Le, 2016) etc. Compared with most compression methods that require huge training costs and access to complete training data, the post-training quantization (PTQ) technique  (Frantar et al., 2022a; Lin et al., 2023; Xiao et al., 2022; Wei et al., 2023) operating with limited calibration data and computational resources is more in demand for compressing LLMs.
Previous PTQ methods (Wei et al., 2023; Xiao et al., 2022; Shao et al., 2023) for LLMs mainly focus on designing accurate outliers suppression schemes for floating-point activations, and then conduct quantization with vanilla policy. They locate the position and optimize the channel scale of activations with outliers in a layer-by-layer or block-by-block manner, as shown in Fig. 1. However, these methods lead to accumulated errors and sub-optimal quantization across the entire model for not taking the dependency across blocks into consideration. Moreover, these PTQ methods suffer from the absence of joint optimization for reducing quantization errors of weights and activations, resulting in significant performance degradation when compressing LLMs into low-bit settings, such as W2A16 and W4A4.
In this paper, we present CBQ, a comprehensive PTQ solution tailored for LLMs, which incorporates outlier suppression, learnable quantization step sizes, and adaptive rounding in a unified cross-block framework, as shown in Fig. 2. CBQ introduces a cross-block dependency (CBD) into block-wise reconstruction, helping mitigate the issue of accumulated errors across the entire model. Our approach ensures that multiple transformer blocks within a sliding window with overlapping are optimized simultaneously. This overlapping window mechanism, complemented by a novel homologous reconstruction scheme, allows for more effective and non-local optimization of quantization parameters. To reduce reconstruction difficulties posed by extreme weight and activation outliers, CBQ incorporates a novel coarse-to-fine preprocessing (CFP) strategy. CFP utilizes a quartile criterion to coarsely determine the range of outliers and then precisely pinpoint the locations of the outliers based on feature distributions. Then we truncate weight outliers and perform equivalent scaling on the activation outliers. Furthermore, CBQ introduces a LoRA-Rounding technique, employing two low-rank matrices to learn the adaptive rounding values for quantized weight, which helps rectify weight quantization errors while preserving training efficiency. It is worth noting that the step sizes of weights and activations, and the rounding matrices are jointly optimized in our proposed quantization framework. CBQ’s innovative approach in cross-block reconstruction, advanced outlier processing, and adaptive rounding establish it as a groundbreaking solution in the field of PTQ for LLMs.
The contributions of this paper are summarized as follows:
  • We propose a unified PTQ method tailored to LLMs based on cross-block reconstruction, which introduces a cross-block dependency mechanism to reduce accumulated errors in a block-wise manner.
  • We design a coarse-to-fine pre-processing strategy to truncate weight outliers and equivalently scale activation outliers to reduce the reconstruction difficulty.
  • We present LoRA-Rounding, an efficient adaptive rounding scheme that optimizes rounding errors of weight quantization via two low-rank matrices.
  • Extensive experiments demonstrate that CBQ achieves low-bit quantization settings such as W4A4, W4A8, and W2A16, and outperforms state-of-the-art methods across various models and benchmark datasets.
2 Related works
@@
The post-training quantization (PTQ) algorithm (Nagel et al., 2021) converts the pre-trained full-precision network into a fixed-point network with a few unlabeled calibration data and computational overhead, which enables fast deployment on various devices. Recent post-training quantization methods have been widely explored in vision models (Liu et al., 2021; Hubara et al., 2021; Frantar & Alistarh, 2022; Cai et al., 2020). Some techniques like AdaQuant (Hubara et al., 2020), AdaRound (Nagel et al., 2020), and BRECQ (Li et al., 2021) minimize the distance between floating point and quantized model outputs to optimize quantization parameters. While BRECQ incorporates Fisher information and jointly optimizes layers within each residual block, it still obtains sub-optimal performance for not capturing interactions across neighboring residual blocks. The proposed CBQ improves quantization accuracy that accounts for dependencies between adjacent blocks.
@@
Existing large language models such as BLOOM (Laurençon et al., 2022), OPT (Zhang et al., 2022), and LLAMA (Touvron et al., ) contain tens of billions of parameters, and require massive memory footprint and computation requirements in the inference. Recent works have been proposed to compress LLMs with post-training quantization methods that do not require a complete training procedure and access to a full training dataset. LLM.int8() (Dettmers et al., ), ZeroQuant (Yao et al., 2022) and nuQmm (Park et al., 2022) focus on quantizing the parameters with mixed-precision decomposition scheme, representing the outliers with 16-bit and others with 8-bit. These methods can not truly accelerate the inference of LLMs for that is hard to implement on hardware. Other methods like GPTQ (Frantar et al., 2022b) and AWQ (Lin et al., 2023) can efficiently quantize LLMs but they focus on FP16 activations and INT4 weights, which can not benefit from the integer matrix multiplication of existing AI accelerators. Additionally, Some methods like SmoothQuant (Xiao et al., 2022), Outlier Suppression (Wei et al., 2022b), Outlier Suppression+ (Wei et al., 2023) and QLLM (Liu et al., 2023) aim at processing activation outliers and lack optimization for the weight quantization. Moreover, these methods rely on hand-craft quantization strategies which are tuned based on extensive experimentation for optimization. Recent block reconstruction-based PTQ method OmniQuant (Shao et al., 2023), suffers from error accumulation from the lack of cross-block dependency. In contrast, CBQ incorporates outlier suppression, learnable step sizes, and adaptive rounding in a unified cross-block framework, which enables efficient and high-performance quantization for LLMs.
3 Methods
In this section, we introduce the proposed cross-block quantization framework tailored to LLMs. As illustrated in Fig. 2, CBQ firstly handles the outliers of weights and activations, and then jointly learns step sizes of weights and activations and weight-rounding matrices in a cross-block manner. CBQ reconstructs the output feature of the last block in each sliding window based on the corresponding supervision of the full-precision model.
We first give a brief introduction to quantization, which aims to represent weights and activations of float model with lower bit-width for reducing the memory and computational cost. Given floating-point weights $W$ and activations $X$, the $k$-bit quantized and de-quantized processing of weights and activations can be represented as
$Q(X)=S_{X}\cdot\left\lfloor\frac{X}{S_{X}}\right\rceil;\ \ \ \ \ Q(W)=S_{W}\cdot(\left\lfloor\frac{W}{S_{W}}\right\rfloor+\Delta_{W}),$ (1)
[…]
Figure 3: Diagram of the homologous reconstruction scheme. $\mathcal{L}_{homo\_rec}$ and $\mathcal{L}_{norm\_rec}$ represent the homologous reconstruction loss and the normal reconstruction loss, respectively.
3.1 Cross-block reconstruction
Most existing post-training quantization methods (Xiao et al., 2022; Lin et al., 2023; Wei et al., 2023; Frantar et al., 2022a) for LLMs conduct calibration in a layer-wise manner, optimizing the quantization step sizes of weights and inputs, respectively. OmniQuant (Shao et al., 2023) adopts a learnable method in a block-wise manner but ignores the issue of accumulated errors across the entire model. Besides, previous quantization methods get the rounding values of weights based on the policy of rounding to the nearest integer value, which is not the best for it not adapts to the data and the task loss (Nagel et al., 2020). In this paper, we propose to jointly optimize the step sizes and weight rounding values ($S_{X}$, $S_{W}$ and $\Delta_{W}$) in a unified cross-block framework. For a large language model with $K$ transformer blocks, the objective for joint optimization of the quantization parameters within $k^{th}$ block is formulated as
$\mathop{\arg\min}\limits_{S^{k}_{X},S^{k}_{W},\Delta^{k}_{W}}\mathbb{E}(T_{k}(W^{k},X^{k}),T_{k}(Q(W^{k}),Q(X^{k})),$ (2)
where $1\leq k\leq K$, $T(\cdot)$ is a transformer block, $W^{k}$ and $X^{k}$ are weights and activations within the $k^{th}$ block, respectively. $S_{X}^{k}$ and $S_{W}^{k}$ are quantization step sizes of activation and weight quantization in the block, respectively, which are learnable during optimization. And $\mathbb{E}(\cdot)$ represents the metric to evaluate the reconstruction errors between outputs of quantized block and full-precision block.
Cross-block dependency.
While block-wise reconstruction is effective in saving computational memory, it only considers the local information within each block, neglecting the dependency across different blocks. The reconstruction error will accumulate as the stack of transformer blocks, incurring suboptimal solutions for quantization parameters To tackle this issue, we introduce a cross-block dependency (CBD) scheme using a sliding window approach. This scheme enables the simultaneous optimization of multiple blocks within the window. Furthermore, the two adjacent sliding windows have overlapping blocks, ensuring that the blocks between the windows are also interconnected. The CBD scheme enhances the connectivity and cooperation between blocks, enabling them to jointly contribute to the quantization process. This holistic optimization strategy leads to better overall performance and addresses the limitations of block-wise reconstruction in capturing cross-block dependencies. We formulate the optimization with the CBD scheme as
where $1\leq i\leq k\leq K$, $T_{i,k}$ represents the blocks from block $i$ to block $k$ within one sliding window, and the same applies to the symbols $S_{X}^{i,k}$, $S_{W}^{i,k}$ and $\Delta_{W}^{i,k}$.
Although the CBD technique helps reduce the reconstruction difficulty, it is important to note that it cannot completely eliminate the accumulated errors and the optimization within each window remains local. To address this limitation and further enhance the reconstruction process while ensuring stability, we introduce an additional homologous reconstruction scheme, as illustrated in Fig. 3. In this scheme, we input the output of the previous block of the quantized model into the current block of the full-precision model. This generates an additional output that is homologous to the outputs of the quantized model. By minimizing the reconstruction errors between these two homologous outputs, CBD could help improve the overall accuracy and stability of the reconstruction process. Additionally, the optimization object is the average of the homologous reconstruction error $\mathcal{L}_{homo\_rec}$ and the normal reconstruction error $\mathcal{L}_{norm\_rec}$:
$\mathcal{L}_{rec}=$ $\underbrace{\mathbb{E}(T_{i,k}(W^{i,k},X^{i,k}),T_{i,k}(Q(W^{i,k}),Q(X^{i,k}))}_{\mathcal{L}_{norm\_rec}}+$ (4)
$\underbrace{\mathbb{E}(T_{i,k}(W^{i,k},Q(X^{i,k})),T_{i,k}(Q(W^{i,k}),Q(X^{i,k}))}_{\mathcal{L}_{homo\_rec}}.$
For the distance metric, we incorporate $\mathcal{L}_{2}$ and Kullback-Leibler divergence (KLD) loss (Kullback & Leibler, 1951) to measure reconstruction error. KLD computes the likelihood distribution between output features that undergo the softmax function. It tends to suppress outliers in the feature space and enhance the robustness of the optimization process. By incorporating both terms, our method captures both the spatial distance and the distribution discrepancy, leading to a more comprehensive and robust optimization process. Then the distance metric is formulated as:
where $h_{1}$ and $h_{2}$ are hidden states from the outputs of full-precision blocks and quantized blocks, respectively. $\sigma$ is the softmax function. $\left||\cdot\right||_{2}$ represents the $\mathcal{L}_{2}$ distance and $D_{KL}(\cdot)$ represents the KLD distance. We provide the ablation study on the loss functions in Table 6.
Input: The input tensor $X$,
The balancing coefficient $\lambda_{1}$, $\lambda_{2}$
Output: Outlier $O$
Coarse-grained Detection;
$X_{\text{sorted}}=\text{Sort}(X)$;
$Q_{1}=X[n/4],Q_{3}=X[3n/4]$;
$IQR=Q_{3}-Q_{1}$;
$T=Q_{3}+\lambda_{1}IQR$;
$O=\{x|x>T,x\in X_{\text{sorted}}\}$;
[…]
$N=\textrm{Len}(O),M^{*}=\textbf{INF}$;
foreach  $i$ = 0 to $N$ do
2       $O_{outlier}=O_{i:N}$; $O_{reserved}=O_{0:i}$; $M_{intra}=\textrm{Var}(O_{reserved})$; $M_{inter}=(\textrm{Min}(O_{outlier})-\textrm{Max}(O_{reserved}))^{2}$; $M=M_{inter}-\lambda_{2}M_{intra}$; if $M>M^{*}$ then
3             $O^{*}=O_{outlier}$; $M^{*}=M$.
4       end if
5
6 end foreach
Algorithm 1 Coarse-to-Fine Preprocessing
3.2 Coarse-to-fine pre-processing

### v5

@@ Abstract
Post-training quantization (PTQ) has played a pivotal role in compressing large language models (LLMs) at ultra-low costs. Although current PTQ methods have achieved promising results by addressing outliers and employing layer- or block-wise loss optimization techniques, they still suffer from significant performance degradation at ultra-low bits precision. To dissect this issue, we conducted an in-depth analysis of quantization errors specific to LLMs and surprisingly discovered that, unlike traditional sources of quantization errors, the growing number of model parameters, combined with the reduction in quantization bits, intensifies inter-layer and intra-layer dependencies, which severely impact quantization accuracy. This finding highlights a critical challenge in quantizing LLMs. To address this, we propose CBQ, a cross-block reconstruction-based PTQ method for LLMs. CBQ leverages a cross-block dependency to establish long-range dependencies across multiple blocks and integrates an adaptive LoRA-Rounding technique to manage intra-layer dependencies. To further enhance performance, CBQ incorporates a coarse-to-fine pre-processing mechanism for processing weights and activations. Extensive experiments show that CBQ achieves superior low-bit quantization (W4A4, W4A8, W2A16) and outperforms existing state-of-the-art methods across various LLMs and datasets. Notably, CBQ only takes 4.3 hours to quantize a weight-only quantization of a 4-bit LLAMA1-65B model, achieving a commendable trade off between performance and efficiency.
1 Introduction
Large language models (LLMs) (Wei et al. (2022a); Radford et al. ; Zhang et al. ; Brown et al. (2020b); Dettmers et al. (2022)), have sparked immense academic and industrial interest owing to their remarkable performance in handling complex natural languages tasks (Hendrycks et al. (2020b); Bisk et al. (2020b); He et al. (2017); Ainslie et al. (2023); Liu et al. (2024b)). During to significant computational resources for inference and deployment, the post-training quantization (PTQ) technique  (Choi et al. (2018); Frantar et al. (2022a); Nagel et al. (2019); Wei et al. (2023); Li et al. (2025)) operating with limited calibration data and computational resources is more in demand for compressing LLMs.
Existing PTQ methods typically optimize models on a layer or block basis, addressing outliers (Wei et al. (2022b; 2023); Chee et al. (2024); Liu et al. (2024a)) and employing first- or second-order optimization techniques (predominantly optimizing models on a layer-by-layer or block-by-block basis) (Shao et al. (2023); Frantar et al. (2022b); Liu et al. (2023a)). However, these approaches often suffer from significant performance degradation, particularly in low-bit settings such as W2A16 and W4A4, as illustrated in Table 1, due to inherent limitations. Previous work, like AdaRound (Nagel et al. (2020)), analyzed rounding errors and showed that simple rounding is not always the optimal quantization strategy, greatly improving quantization for CNNs. This inspired us to analyze quantization loss for LLMs, comparing high-bit and low-bit scenarios. We found that in low-bit quantization, intra-layer and inter-layer dependencies within models become more pronounced, especially as model size increases. This indicates that previous methods, whether focused on optimizing quantization parameters within a layer or block through first- or second-order techniques, or on refining rounding errors, fall short of achieving optimal outcomes. Instead, it is essential to fully account for the inter-layer and intra-layer relationships.
To address this, we propose CBQ, a cross-block reconstruction-based PTQ method tailored for LLMs, surpassing traditional layer-wise and block-wise reconstruction techniques. CBQ introduces a cross-block dependency (CBD) into block-wise reconstruction, maintaining the integrity of the model’s internal dependencies during quantization. Our approach optimizes multiple transformer blocks within a sliding window with overlapping, allowing for more effective and non-local optimization of quantization parameters. Using the CBD method, CBQ incorporates a LoRA-Rounding technique, employing two low-rank matrices to learn adaptive compensation values for quantized weights. Notably, we jointly optimize the compensation matrices and the step sizes of weights and activations within the overlapping window, which helps manage intra-layer dependencies to rectify weight quantization errors while preserving training efficiency. Furthermore, CBQ introduces a novel unified coarse-to-fine pre-processing (CFP) strategy from a statistical perspective to evaluate outliers in weights and activations, precisely handling outliers while minimizing damage to normal channels. CFP employs a quartile criterion to initially estimate the range of outliers and then assesses the intra-class and inter-class distances between outliers and normal values to precisely identify their locations. This approach facilitates the truncation of weight outliers and the application of equivalent scaling to activation outliers.
The contributions of this paper are summarized as follows:
  • We performed a comprehensive analysis of the error sources in low-bit quantization scenarios for LLMs, and theoretically demonstrated the significant impact of intra-layer and inter-layer dependencies on the effectiveness of model quantization.
  • We propose CBQ, a unified PTQ method designed for LLMs, incorporating a cross-block reconstruction strategy that introduces a Cross-Block Dependency (CBD) mechanism to preserve the model’s internal dependencies during quantization, and LoRA-Rounding to utilize intra-layer dependencies for optimizing adaptive compensation matrices.
  • We design a coarse-to-fine pre-processing strategy (CFP) that can simultaneously detect and manage outliers in both weights and activations, effectively preventing disruption to normal activation channels and weights.
  • Extensive experiments demonstrate the effectiveness of our method in ultra-low bit quantization settings such as W4A4, W4A8, and W2A16. Notably, it outperforms state-of-the-art methods across diverse models and benchmark datasets.
2 Motivation
Figure 1: (a) Visualization of the absolute values of the Hessian matrix for weights within a single layer of LLAMA-7B, (b) Hessian matrix visualization of the loss with respect to the scale across 32 layers of LLAMA-7B, and (c) the relationship between the average scale of the first two transformer blocks in LLAMA-7B and the corresponding loss.
To analyze the sources of quantization errors in large models when quantizing weights or activations, we assume a matrix $M$ representing a set of weights or activations as the current quantization target, and $\mathcal{L}$ denotes the quantization loss of the model under this matrix. Let $\varepsilon$ denote a small perturbation introduced by quantization and $\mathcal{L}(M)$ represent the task loss that we aim to minimize. Then,we can derive the following equation within the Taylor expansion:
$\mathbb{E}[\mathcal{L}(M+\varepsilon)-\mathcal{L}(M)]\approx\mathbb{E}[\varepsilon^{T}\cdot\frac{\partial\mathcal{L}}{\partial M}+\frac{1}{2}\varepsilon^{T}\frac{\partial^{2}\mathcal{L}}{\partial M^{2}}\varepsilon+O(||\varepsilon||^{3})]\approx\varepsilon^{T}\cdot g^{(M)}+\frac{1}{2}\varepsilon^{T}\cdot\textbf{H}^{(M)}\cdot\varepsilon$ (1)
As discussed in previous work  (Frantar et al. (2022b)), The error $\varepsilon$ introduced by quantization is sufficiently small, the higher-order terms in the Taylor expansion can be neglected. Therefore, we analyze the first- and second-order terms, $g^{(M)}$ and $\textbf{H}^{(M)}$, which can be defined as follows.
$g^{(M)}=\mathbb{E}[\nabla_{M}\mathcal{L}(M)]=\sum_{i}^{K}\frac{\partial\mathcal{L}}{\partial M_{i}}$ (2)
$\textbf{H}^{(M)}=\mathbb{E}[\nabla_{M}^{2}\mathcal{L}(M)]=\sum_{i}^{K}\sum_{j}^{K}\frac{\partial^{2}\mathcal{L}}{\partial M_{i}\partial M_{j}}$ (3)
Let $K$ denote the number of elements in the LLM involved in the quantization. From equations 123, it can be observed that when the quantization perturbation $\varepsilon$ is small, $||\varepsilon||^{2}$ is also small, allowing us to disregard the implications of the equation 3. In this case, the quantization error is primarily related to the current quantization target $M$, analogous to high-bit quantization. However, when performing low-bit quantization, $||\varepsilon||^{2}$ increases, necessitating consideration of the impact described by Equation 3. This indicates that when $i\neq j$, relationships between different $M$ are introduced. This relationship manifests in two aspects: when quantizing a single layer, it reflects intra-layer dependencies among parameters, and when quantizing the entire model, inter-layer dependencies must also be considered. Furthermore, given that the complexity of the Hessian matrix H is proportional to $\mathcal{O}(n^{2})$, where $n$ represents the number of parameters, the growth in model size, both in terms of parameters and layers, leads to a marked intensification of intra-layer and inter-layer dependencies.
To better illustrate intra-layer and inter-layer dependencies, we visualize Equation 3 for both individual layers and the entire model using LLAMA-7B. Additionally, we present visualizations of the dependencies between adjacent blocks, as referenced in Figure 1.
By analyzing Figure 1, we observe a notable increase in the values of off-diagonal elements during lower-bit quantization. This increase indicates a strengthening of both inter-layer and intra-layer dependencies, with closer elements exhibiting stronger correlations. Furthermore, comparisons of the scales between adjacent layers provide a clearer understanding of the substantial impact that inter-layer dependencies have on final quantization outcomes in low-bit scenarios.
Therefore, taking into account both intra-layer and inter-layer dependencies, we present the quantization framework for LLMs under low-bit settings, which can be expressed by the following equation:
$\mathop{\arg\min_{h\subseteq\textbf{H}_{+}}\sum_{k\in h}}\mathbb{E}(T_{k}(W^{k},X^{k}),QT_{k}(Q(W^{k})+\Delta_{W}^{k},Q(X^{k})),$ (4)
where $T$ and $QT$ represent the floating-point and quantized transformer blocks, respectively. $Q_{\cdot}(\cdot)$ represents the quantization process. $\mathbb{E}(\cdot)$ represents the metric to evaluate the reconstruction errors between outputs of quantized block and full-precision block. We jointly optimize all transformer blocks with inter-layer dependencies while compensating for intra-layer relationships using $\left\{\Delta_{W}^{k}|k\subseteq\textbf{H}_{+}\right\}$.
3 Method
@@
In this section, we introduce the proposed cross-block quantization framework tailored to LLMs. As illustrated in Fig. 2, CBQ firstly handles the outliers of weights and activations, and then jointly learns step sizes of weights and activations and weight-compensation matrices in a cross-block manner. CBQ reconstructs the output feature of the last block in each sliding window based on the corresponding supervision of the full-precision model.
3.1 Cross-block reconstruction
To maintain inter-layer dependencies, it is necessary to optimize the layers with significant dependencies together. As shown in Figure 1, the strongest dependencies are typically observed between adjacent layers. Therefore, we introduce a cross-block dependency (CBD) scheme using a sliding window approach. This scheme enables the simultaneous optimization of multiple blocks within the window. Furthermore, the two adjacent sliding windows have overlapping blocks, ensuring that the blocks between the windows are also interconnected. The CBD scheme enhances the connectivity and cooperation between blocks, enabling them to jointly contribute to the quantization process. This holistic optimization strategy leads to better overall performance and addresses the limitations of block-wise reconstruction in capturing cross-block dependencies. We formulate the optimization with the CBD scheme as
where $1\leq i\leq k\leq K$, $T_{i,k}$ represents the blocks from block $i$ to block $k$ within one sliding window, and the same applies to the symbols $S_{X}^{i,k}$, $S_{W}^{i,k}$ and $\Delta_{W}^{i,k}$. The optimization object $\mathcal{L}_{rec}$ is as follow:
$\mathcal{L}_{rec}=\mathbb{E}(T_{i,k}(W^{i,k},X^{i,k}),T_{i,k}(Q(W^{i,k}),Q(X^{i,k}))$ (6)
For the distance metric, we incorporate $\mathcal{L}_{2}$ and Kullback-Leibler divergence (KLD) loss (Kullback & Leibler (1951)) to measure reconstruction error. KLD computes the likelihood distribution between output features that undergo the softmax function. It tends to suppress outliers in the feature space and enhance the robustness of the optimization process. By incorporating both terms, our method captures both the spatial distance and the distribution discrepancy, leading to a more comprehensive and robust optimization process. Then the distance metric is formulated as:
where $h_{1}$ and $h_{2}$ are hidden states from the outputs of full-precision blocks and quantized blocks, respectively. $\sigma$ is the softmax function. $\left||\cdot\right||_{2}$ represents the $\mathcal{L}_{2}$ distance and $D_{KL}(\cdot)$ represents the KLD distance. We provide the ablation study on the loss functions in Appendix B.Table 5.
3.2 LoRA-Rounding for weight quantization
AdaRound (Nagel et al. (2020)) introduces to learn a better weight-rounding matrix for post-training quantization that adapts to the data and the task loss. As shown in Eq. 8, we can obtain the weight-rounding matrix $\Delta_{W}\in\mathbb{R}^{d\times k}$ with a learnable matrix $V\in\mathbb{R}^{d\times k}$ with a rectified sigmoid function:
When the transformer blocks are within the overlap of the CBD sliding window mechanism, the rounding matrix can serve as an effective representation of intra-layer dependencies. We utilize it as a compensation matrix and jointly optimize it with the quantization step sizes for weights and activations, which can be expressed as follows:
$\mathop{\arg\min}\limits_{S_{X}^{i,k},S_{W}^{i,k},\Delta_{W}^{j,k}}\mathbb{E}(T_{i,k}(W^{i,k},X^{i,k}),T_{i,k}(Q(W^{i,k})+\Delta_{W}^{j,k},Q(X^{i,k}))$ (9)
[…]
However, as shown in Experiment Table 3(c), we found that LLMs with billion-level parameters result in an exceptionally large $\Delta_{W}^{j,k}$, which can lead to significant computational overhead and substantially impact the convergence of training. (Shao et al. (2023)) has also mentioned that AdaRound cannot be applied to models with billions of parameters due to the vast solution space, which aligns with our experimental findings. Thus, we employ low-rank adaptive learning on the compensation matrices, decomposing $V$ with much smaller low-rank matrices, and only optimize them in post-training quantization, the decomposition is defined as:
Where the rank $r<<\textrm{min}(d,k)$, we utilize a random Gaussian initialization for $A_{1}$ and zero for $A_{2}$, thus $\Delta_{W}$ is set to zero at the beginning of post-training quantization. During training, each element of $V$ is encouraged into 0 or 1 with a regularizer loss:
$\mathcal{L}_{com}=\sum_{i,j}1-|2\Delta_{W}{(i,j)}-1|^{\beta},$ (12)
Where $\beta$ is a annealing factor. Following  (Nagel et al. (2020)), $\beta$ is set to higher in the initial phase and set to lower in the later phase of the optimization to encourage it to converge to 0 or 1. We also conduct $\Delta_{W}=\left\lfloor\Delta_{W}\right\rceil$ in the later phase of the optimization to force each element into $\{0,1\}$ exactly.
Compared with vanilla AdaRound for LLMs. The proposed LoRA-Rounding reduces the number of learnable parameters from $d\times k$ to $(d+k)\times r$ and changes the training strategy, significantly accelerating the optimization process, we conduct ablation experiments in the next section 5.3.
3.3 Overall loss
In summary, by leveraging CBD and a low-rank decomposition of the weight-compensated matrix, We slide the window to the last block with an interval and update all the quantization parameters $S_{W},S_{X},A_{1},A_{2}$ within the window, ensuring the preservation of both intra-layer and inter-layer relationships of the model, thereby achieving optimal performance. The total loss for optimizing the $i^{th}$ block to the $k^{th}$ block within a sliding window is formulated as
$\mathcal{L}_{total}=\mathcal{L}_{rec}+\gamma\mathcal{L}_{com},$ (13)
where the $\gamma$ is the hyper-parameter to balance the reconstruction error and compensation error.
3.4 Coarse-to-fine pre-processing
Outlier handling is crucial in quantizing LLMs. Figure 3 in Appendix F illustrates the prevalent outliers in weights and activations, which pose significant challenges to the quantization process. Although there are many existing studies based on outliers problem, these studies typically focus on outliers in either weights or activations individually, such as in (Chee et al. (2024); Wei et al. (2022b); Xiao et al. (2022)). However, there is no precise strategy that can simultaneously detect and handle outliers in both weights and activations. This single-mode approach can potentially damage normal activation channels and weights due to incorrect outlier detection. To address this issue, we discard the previous assumption of normal distributions for weights and activations (Wu et al. (2023)), and based on statistical principles (Massart et al. (2005)), propose a coarse-to-fine pre-processing strategy to decouple the outlier handling in activations and weights. Relevant theoretical details can be found in Appendix F.
The comprehensive algorithm of the outlier detection is illustrated in Algorithm 1 in Appendix K, which is divided into two stages.
Coarse-grained detection.
In the first stage, we perform coarse-grained detection by calculating the lower and upper quartile values ($Q_{1}$ and $Q_{3}$) and the interquartile range ($IQR$) (Massart et al. (2005)) in the numerical distribution (either activations or weights). Based on these calculations, we obtain a coarse outlier set $O=\{x|x>T,x\in X\}$, where $T=Q_{3}+\lambda_{1}IQR$ and $\lambda_{1}$ is set to 1.5. This stage greatly reduces the search space for outlier detection.
Fine-grained detection.
Removing outliers in weights has minimal impact on performance, whereas outliers in activations, particularly in specific channels, can greatly affect performance if directly removed. Consequently, our approach involves truncating weight outliers and scaling outliers in activations based on the detected outliers in both weights and activations. Figure 3 in Appendix F provides visual evidence of weight outliers being truncated within the outlier group.
$s_{i}=\sqrt{\textrm{Max}(|X_{i}|)/\textrm{Max}(O^{*})}.$ (14)
This scaling factor is then applied to update weights and activations following prior work (Wei et al. (2023)) to counteract destabilizing fluctuations from remaining outliers.
4 Related Work
@@
The post-training quantization (PTQ) algorithm (Nagel et al. (2021); Wu et al. (2020; 2023); Zhang et al. (2018)) converts the pre-trained full-precision network into a fixed-point network with a few unlabeled calibration data and computational overhead, which enables fast deployment on various devices. Recent post-training quantization methods have been widely explored in vision models (Liu et al. (2021); Hubara et al. (2021); Frantar & Alistarh (2022); Cai et al. (2020); Li et al. (2022)). Some techniques like AdaQuant (Hubara et al. (2020)), AdaRound (Nagel et al. (2020)), and BRECQ (Li et al. (2021)) minimize the distance between floating point and quantized model outputs to optimize quantization parameters. While BRECQ incorporates Fisher information and jointly optimizes layers within each residual block, it still obtains sub-optimal performance for not capturing interactions across neighbouring residual blocks. The proposed CBQ improves quantization accuracy that accounts for dependencies between adjacent blocks.
@@
Existing large language models such as BLOOM (Laurençon et al. (2022)), OPT (Zhang et al. (2022)), and LLAMA (Touvron et al. ; Touvron et al. (2023)) contain tens of billions of parameters, and require massive memory footprint and computation requirements in the inference (Ashkboos et al. (2023); Wang et al. (2010); Bolya & Hoffman (2023); Brown et al. (2020a); Jacob et al. (2018)). Recent works have been proposed to compress LLMs with post-training quantization methods that do not require a complete training procedure and access to a full training dataset. LLM.int8() (Dettmers et al. ), ZeroQuant (Yao et al. (2022)) and nuQmm (Park et al. (2022)) focus on quantizing the parameters with mixed-precision decomposition scheme, representing the outliers with 16-bit and others with 8-bit. These methods can not truly accelerate the inference of LLMs for that is hard to implement on hardware. Other methods like GPTQ (Frantar et al. (2022b)) and AWQ (Lin et al. (2023)) can efficiently quantize LLMs but they focus on FP16 activations and INT4 weights, which can not benefit from the integer matrix multiplication of existing AI accelerators. Additionally, Some methods like SmoothQuant (Xiao et al. (2022)), Outlier Suppression (Wei et al. (2022b)), Outlier Suppression+ (Wei et al. (2023)) and QLLM (Liu et al. (2023a)) aim at processing activation outliers (Zhao et al. (2019)) and lack optimization for the weight quantization. Moreover, these methods rely on hand-craft quantization strategies which are tuned based on extensive experimentation for optimization. Recent block reconstruction-based PTQ method OmniQuant (Shao et al. (2023)),QLLM (Liu et al. (2023a)), have experienced significant accuracy degradation in low-bit settings. In contrast, CBQ introduces a more precise outlier detection strategy and optimizes the reconstruction process through CBD and LoRA-Rounding mechanisms by maintaining both intra-layer and inter-layer dependencies.
5 Experiments
Table 1: Evaluation on multiple zero-shot datasets with the accuracy $\uparrow$ metric, where the Mutual dataset is evaluated with the Mean Reciprocal Rank/Recall@$1$/Recall@$2$ metrics. CBQ represents that the experiments conducted with 2-bit weight-only quantization did not fully quantize the model but only the FC2 layer of the first and last transformer blocks are converted to 4-bit precision.